# Survival Modeling from Whole Slide Images via Patch-Level Graph Clustering and Mixture Density Experts

### v1

@@ 4. Feature Integration: We first concatenate refined patch features from all clusters as follows:
$R^{\prime}_{\mathrm{expanded}}=$ $\mathrm{mean}(R^{\prime},\mathrm{dim}=1,\ \mathrm{keepdim=True})$
.$\mathrm{.expand}(-1,\ \widetilde{\mathrm{P}}.\mathrm{shape}[1],\ -1)$ (19)
@@ 1. Transformed Probability Density Function (TPDF):
$\mathrm{TPDF}(t\mid\mathrm{z}_{\mathrm{WSI}},e)=$ $\left|\frac{dy}{dt}\right|\cdot\sum_{i=1}^{K}\lambda_{i}^{(e)}(\mathrm{z}_{\mathrm{WSI}})\cdot$
@@ 4.5 Training Objective
$\mathrm{L}_{\mathrm{NLL}}=$ $-c\cdot\log(\mathrm{TPDF}(t_{d}\mid\mathrm{z}_{\mathrm{WSI}}))$
@@ 5.3 Performance Metrics
To comprehensively assess model performance beyond traditional metrics, we adopted enhanced evaluation measures that capture both discrimination and calibration over time. While the standard concordance index (C-Index) [16] provides a global ranking of predicted risks, it is limited in temporal granularity. Therefore, we employed the time-dependent concordance (TDC), which assesses the proportion of correctly ranked patient pairs at multiple time points within a predefined interval $[0,\tau]$, offering a dynamic perspective on discriminative ability. In addition, we used the Brier Score (BS) to evaluate the calibration of predicted survival probabilities by measuring the mean squared error between predicted and observed outcomes. To account for prediction accuracy across the entire time horizon, we calculated the Integrated Brier Score (IBS), which integrates the BS over the interval $[0,\tau]$. Higher TDC and lower IBS values indicate superior model performance. All reported results are presented as mean ± standard deviation across validation folds.

### v2

@@ 4. Feature Integration: We first concatenate refined patch features from all clusters as follows:
$R^{\prime}_{\mathrm{expanded}}=\$$\mathrm{mean}(R^{\prime},\mathrm{dim}=1,\ \mathrm{keepdim=True})$
$\mathrm{.expand}(-1,\ \widetilde{\mathrm{P}}.\mathrm{shape}[1],\ -1)$ (19)
@@ 1. Transformed Probability Density Function (TPDF):
$\mathrm{TPDF}(t\mid\mathrm{z}_{\mathrm{WSI}},e)=\$$\left|\frac{dy}{dt}\right|\cdot\sum_{i=1}^{K}\lambda_{i}^{(e)}(\mathrm{z}_{\mathrm{WSI}})\cdot$
@@ 4.5 Training Objective
$\mathrm{L}_{\mathrm{NLL}}=\$$-c\cdot\log(\mathrm{TPDF}(t_{d}\mid\mathrm{z}_{\mathrm{WSI}}))$
@@ 5.3 Performance Metrics
To comprehensively assess model performance beyond traditional metrics, we adopted enhanced evaluation measures that capture both discrimination and calibration over time. While the standard concordance index (C-Index) [16] provides a global ranking of predicted risks, it is limited in temporal granularity. Therefore, we employed the time-dependent concordance (TDC), which assesses the proportion of correctly ranked patient pairs at multiple time points within a predefined interval $0,\tau]$, offering a dynamic perspective on discriminative ability. In addition, we used the Brier Score (BS) to evaluate the calibration of predicted survival probabilities by measuring the mean squared error between predicted and observed outcomes. To account for prediction accuracy across the entire time horizon, we calculated the Integrated Brier Score (IBS), which integrates the BS over the interval $0,\tau]$. Higher TDC and lower IBS values indicate superior model performance. All reported results are presented as mean ± standard deviation across validation folds.